OpenStudy (breanna77):

The midpoint of UV is (5, -11). The coordinates of one point are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (breanna77):

Will give a medal

OpenStudy (bradely):

u(3,5)v(x,y) Mid point (5,-11) (3+x)/2=5 3+x=10 x=7 (y+5)/2=-11 y+5=-22 y=-27 v(7,-27) Step by step answer http://www.mathskey.com/question2answer/

OpenStudy (breanna77):

Oh my gosh @bradely, thank you so much!!